OpenStudy (anonymous):

Find the LCM of y^3-y^2 and y^2-y^2

OpenStudy (zzr0ck3r):

\(y^2-y^2=0\) so \(Lcm(y^2-y^2,y^3-y^2)=Lcm(0,y^3-y^2)\) is undefined, else \(0\) would divide \(Lcm(0,y^3-y^2)\) and of course this cant happen.